OpenStudy (anonymous):

how to solve 6g^3-24g^2+24g

OpenStudy (anonymous):

Factorize it: 6g(g^2 - 4g + 4) which will be 6g . (g-2)^2

OpenStudy (anonymous):

ok how did you get to 6g(g-2)^2 after factorizing?

OpenStudy (anonymous):

By using *** (a-b)^2 = a^2 + b^2 - 2ab *** so g^2 - 4g + 4 = (g-2)^2